OpenStudy (anonymous):

Factor completely: 5x^2+30x+45

OpenStudy (anonymous):

i think it is 2 numbers that produce 45 and add up to 30 a*b=45 a+b=30

OpenStudy (anonymous):

5 (x+3)^2 well its actually this then

OpenStudy (anonymous):

5(x^2 + 6x + 9) 5(x+3)^2

OpenStudy (anonymous):

im sorry im still confused about this. Can you explain I really want to understand this

OpenStudy (anonymous):

you start off by dividing 5 from the whole equation. it's so that you can factor easier. after factoring out the 5 you get a less complex equation. With this you can easily find that 3 times 3 is 9 and 3+3 is 6. with that you can get the answer. 5(x^2 + 6x + 9) you repeat the 3 in both parentheses so you just have only one and put a ^2 next to it. 5(x+3)^2 sorry if it made you even more confused. best of luck to you.

OpenStudy (anonymous):

@timo86m first post. This is false, you have to multiply the constant by 5 before you can use the grouping method. That's why it's best to simplify the equation first by taking out 5 as a common factor.

OpenStudy (anonymous):

@Iamapineapple okay, so when im done multiplying i get 5x^2+30x+225 what's the next thing

OpenStudy (anonymous):

@ineedyouubiebs, using that method probably isn't the best way to do it - factoring 225 in your head isn't easy. And when I said multiply the constant by 5, I didn't actually mean in the equation, just in your head :P I'll explain the easy method: You equation: 5x^2+30x+45

OpenStudy (anonymous):

Take a Common Factor of 5 OUT of the equation, so that means I divide everything by 5 and take it outisde. You should ALWAYS take out a common factor if you can 5(x^2 + 6x + 9) Now I can use the method everyone mentioned. I have to find two number that multiply together to give 9 and add to give 6. So a x b = 9 a x b = 6 I know that the only things that multiply to give 9 are 1x9 or 3x3. 1+9 =10, therefore a and b must be 3 so after you find your numbers, you can write it like this 5[(x+3)(x+3)] Because (x+3) is multiplied twice, I can use something called perfect squares. This means I can write (x+3)(x+3) as (x+3)^2 So your answer it 5(x+3)^2

OpenStudy (anonymous):

OH MY GOD THIS MAKES SENCE NOW, THANK YOU SO MUCH. WOULD YOU MIND WORKING ON ANOTHER EXAMPLE IN ANOTHER POST PLEASE?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yes so i think if you have the standard quadtratic equation ax^2+bx+c then if possible factor out the a first.